# Towards Facilitated Fairness Assessment of AI-based Skin Lesion Classifiers Through GenAI-based Image Synthesis

### v3

@@ 1 Introduction
Melanoma is the deadliest form of skin cancer, with an estimated 325,000 new cases (174,000 men and 151,000 women) and 57,000 deaths (32,000 men and 25,000 women) reported worldwide in 2020 [3]. By 2040, incidence is projected to increase by nearly 50% to 510,000 new cases, and mortality by 68% to 96,000 deaths. Early detection, however, dramatically improves survival. For example, a population-wide screening project in Schleswig-Holstein, Germany, reported a reduction in melanoma mortality of 47% for men and 49% for women [26]. Similarly, screening efforts in Australia decreased melanoma deaths by 59% [58]. These findings underline a crucial insight: despite rising incidence, early and accessible screening can substantially reduce mortality.
Recent advances in Artificial Intelligence (AI), together with the widespread availability of high-resolution smartphone cameras, have created new opportunities for accessible melanoma pre-screening outside clinical environments. AI-assisted diagnostic systems can support clinicians during routine examinations and also provide human-understandable explanations for their decisions [33]. In principle, such systems could even enable remote pre-screening, allowing individuals to capture dermoscopic images themselves and receive preliminary risk assessments [9, 45]. These innovations have considerable societal potential, including earlier detection, reduced diagnostic delays, lower healthcare costs, and decreased clinician workload.
Figure 1: Real melanoma test data (left) such as MILK10k [36] shows limited and imbalanced demographic coverage, while synthetic data (right) can provide complete combinations with balanced sampling, enabling more reliable fairness assessment.
At the same time, the deployment of AI in medicine requires a high level of trustworthiness, especially in light of regulatory developments such as the EU AI Act [14]. Fairness and transparency in both datasets and trained models are essential conditions for reliable clinical use [5, 2, 1, 6, 39, 12]. Fairness assessments depend on access to demographically well-characterized data. The International Skin Imaging Collaboration (ISIC) archive [18] provides one of the most comprehensive publicly available skin lesion datasets [18, 11, 56, 10, 48, 21], including metadata such as sex, age, and Fitzpatrick skin type [51]. Despite its scale and quality, demographic imbalance remains a major limitation and complicates fairness evaluations.
@@
DeepGuide ([35]) DenseNet Derm7pt 0.788 (AUC)
MelaNet ([61]) VGG-GAP ISIC 2016 0.811 (AUROC)
Patch-Lesion ([15]) Dense121 HAM10000 0.855 (F1-Score)
SkinLesionDensenet ([40]) DenseNet201 ISIC 0.878 (F1-Score)
EFFNet ([34]) EfficientNetV2 HAM10000 0.932 (F1-Score)
Sabir and Mehmood [50] EfficientNet-B0 HAM10000 0.970 (F1-Score)
@@
Meanwhile, recent progress in Generative AI (GenAI), particularly diffusion models [22, 47], has enabled high-fidelity and attribute-controlled image synthesis. In this work, we explore whether these advances can support fairness evaluation in melanoma detection. Using the ISIC dataset and LightningDiT [59], a latent diffusion transformer [41], we train a generative model capable of synthesizing realistic dermoscopic images conditioned on key patient attributes. We then examine whether such synthetic cohorts can serve as a reliable resource for assessing algorithmic fairness in publicly available melanoma classifiers. In summary, our work makes the following key contributions.
@@ 2.1 Generative Models for Skin Lesions
Synthetic data generation is widely used to address data scarcity, class imbalance, and privacy concerns in skin lesion datasets. Early approaches based on Generative Adversarial Networks (GANs) [17], such as MelanoGANs [4], DCGAN [43], and LAPGAN [13], demonstrated that realistic dermoscopic images can be synthesized even from limited data. DermGAN [16] and StyleGAN-based methods [7, 24] introduced finer control over lesion characteristics and improved downstream classifier performance. Later work, including StyleGAN2 [25] and federated setups [8], further improved stability, fidelity, and privacy. Diffusion models now dominate high-fidelity medical image synthesis. Stable Diffusion [47] and DreamBooth [49], as applied in Cancer-Net SCa-Synth [54], support few-shot personalization and balanced dataset construction. Extensions of latent diffusion models have also improved control over skin type variation and helped address fairness gaps [57]. Our work differs by focusing on fairness evaluation rather than augmentation. We train a state-of-the-art latent diffusion transformer based on LightningDiT [59] to generate attribute-controlled dermoscopic image cohorts. The model leverages all available ISIC metadata and allows systematic control over demographic and clinical attributes for constructing balanced intersectional test sets.
@@ 2.2 Melanoma Detection Models
Prior research on melanoma classification has employed a variety of datasets, including ISIC, Derm7pt [27], and HAM [56], the latter containing 10,015 dermoscopic images covering seven lesion types. These datasets differ in composition, imaging style, and demographic distribution, which complicates direct model comparison. A further challenge arises from the heterogeneous evaluation metrics reported in the literature. Some studies rely on the F1-score, whereas others report AUC or AUROC, making cross-model comparison nontrivial. Pretrained model availability also varies. Public weights exist for DeepGuide [35], MelaNet [61], and SkinLesionDensenet [40], but not for Patch-Lesion [15] or EFFNet [34]. In this work, we use publicly available pretrained models to ensure reproducibility and consistent evaluation.
@@ 2.3 Fairness in Skin Lesion Classification
Fairness evaluation has gained increasing attention due to persistent performance disparities across demographic subgroups. In [2], the authors proposed an augmentation strategy that combines a source skin tone image with a diagnostic image to mitigate imbalance and skin tone bias. Other approaches incorporate fairness directly into model design. BiaslessNAS [52] integrates fairness constraints throughout the neural architecture search process and reports improvements of 2.55% in accuracy and 65.50% in fairness over baseline models. Compared to previous augmentation-based fairness approaches, our method generates attribute-controlled synthetic images through text prompts. We evaluate fairness in existing pretrained melanoma classifiers [35, 61, 40] using balanced synthetic cohorts, allowing us to investigate the potential of prompt-driven generative models as a tool for systematic fairness assessment.
@@ 3.1 Synthetic Test Data Generation
Figure 2: Overall pipeline of our fairness testing workflow. We first train a generative model based on LightningDiT [59] using the ISIC dataset. Synthetic melanoma test images are then generated via systematic prompt templates. Finally, the synthetic images are used to assess fairness as measured via TPR difference in pre-trained melanoma detection models.
Data: Only publicly available data are used and no new patient information is collected. All dermoscopic images from the ISIC archive (2016–2024) are processed at a resolution of $256\times 256$ and encoded into the latent space using the VA-VAE, resulting in roughly 500,000 latent-image pairs. Each image is paired with a text description constructed from the available annotations, including sex, age group, lesion size, Fitzpatrick skin type, and diagnosis. Data with insufficient attributes are excluded. Text embeddings are produced using CLIP [44], a contrastively trained vision-language model that embeds images and text into a shared latent space and provides strong semantic grounding for conditioning.
@@ 3.3 Fairness Assessment
Metric Definition: Since our synthetic test set contains only melanoma (i.e. positive) cases, we quantify subgroup disparities using TPR disparity [19], defined as the difference between the maximum and minimum true positive rates across PII groups. In a positive-only setting, the proportion of predicted positives in each group coincides with sensitivity. For a single decision threshold $\tau$ shared across all groups:
@@
We compute $\Delta_{\mathrm{TPR}}$ on synthetic cohorts stratified by sex, age group, and Fitzpatrick skin type. To generate a strictly positive-only test set, we use “melanoma” as the disease value in our systematic prompt configuration. The resulting measure, $\Delta_{\mathrm{TPR}}$, therefore provides a meaningful and statistically robust indicator of whether a model systematically favors or disadvantages particular demographic groups [53, 31, 29].
@@
Models: We evaluate three publicly available pretrained melanoma classifiers: DeepGuide [35], MelaNet [61], and SkinLesionDensenet [40]. These models were selected due to the availability of pretrained weights and their relevance in prior benchmark studies. MelaNet and SkinLesionDensenet are trained on ISIC, whereas DeepGuide is trained on the HAM10000 (HAM) dataset, providing complementary training distributions. Architecturally, DeepGuide and SkinLesionDensenet rely on DenseNet backbones, while MelaNet is based on a VGG-GAP design. This diversity in both training data and network architecture offers a representative foundation for examining fairness behavior across classifier families.
@@ 4.1 Generated Synthetic Skin Lesion Images
While the generated images are highly realistic, one limitation is immediately noticeable. The generated images do not became increasingly darker with higher Fitzpatrick skin type. For example, the generated images do not become progressively darker with higher Fitzpatrick skin types. Although this might be the intuitive expectation, prior work [55, 23] shows that lighting, camera settings, and lesion visibility introduce substantial variability, making pixel intensity an unreliable indicator. Further details and visual comparisons with real data exhibiting the same issue are provided in the appendix.
@@ 4.2 Evaluation of Fairness Performance
Effect of Test Set Size: Figure 5 shows cumulative TPR as a function of the number of test images. The data count corresponds to the number of images used and illustrates how recognition accuracy changes as more test data are included. As a rule of thumb, as the number of test images increases, convergence to a stable accuracy is expected [46]. Therefore, if accuracy trends for real and synthetic images converge, fairness simulations can be reliably performed using generated images. For sex, both male and female curves exhibit this convergence behavior, and synthetic testing closely tracks real-image performance as the available data volume increases. In contrast, the age-related curves for real and synthetic data do not fully converge, suggesting that the generator and real data distributions differ more strongly along the age dimension. Age is a factor with substantial intra-group variability (e.g., early versus late decades), which may contribute to these discrepancies and should be examined in future work.
@@ 5.1 Enhancing Dataset Use for Training
In this work, we generated synthetic image test data by training a custom LightningDiT [59]. According to Figure 3, the images show high realism, and Figure 5 confirms a similar trend in the cumulative TPR comparison between real and synthetic images. However, according to Figure 5, age TPR conversion differs between real and synthetic images. This may be due to the training procedure, which determines the precision of the label used for training. The next step to improve the synthetic image generation model is to include a detail attribute in the training procedure of our generative model. Another direction is to add multiple datasets and use different image capturing devices to make the dataset and thus generation process more robust.
@@ 5.2 Quality Control of the Synthetic Images
In our work, we evaluate fairness performance through 81,600 synthetic images generated in a pattern presented in Section 3.2. To ensure the quality of the generated images, further quality control models could be used that evaluate the reliability of the images [30]. The approach may, for example, be to implement quality-judging models for each age, sex, and skin color. Also, due to the significant recent growth of visual language models (VLMs) [60], VLM-as-a-judge may be an approach to perform quality evaluation before model prediction. Hence, the next step in the current architecture is also to consider new models for quality assurance of the test images.
@@ 5.3 Understanding Causes of Unfairness
Our results verified that synthetic image generation can serve as test data for evaluating the fairness of pre-trained models. For instance, when examining the performance of pre-trained models, we confirm that for sex and skin type, MelaNet shows the most fair behavior, with $\Delta_{\mathrm{TPR}}$ of 0.0115 and 0.1322. For age SkinLesionDensnet performs the best with 0.0682. Although we have managed to measure $\Delta_{\mathrm{TPR}}$ for model fairness comparison, we still lack a method for discovering the causes or factors of unfairness. Hence, the next step is to apply concept-based fairness [37, 42, 55] evaluation to understand the actual cause of the fairness issue.
@@ 5.4 Various Diagnoses for Testing
We focused on TPR assessed through the disparity in true positive rates ($\Delta_{\mathrm{TPR}}$) under a melanoma-only (positive) test design. To broaden the scope of fairness evaluation, we aim to extend the generator and evaluation protocol to multiple diagnostic categories (e.g., benign nevi and other lesion types) and to multi-class classifiers in the future. This will enable (i) analysis of misclassification patterns (false positives and false negatives) by PII group, (ii) computation of additional group fairness criteria beyond $\Delta_{\mathrm{TPR}}$, such as equalized odds [19], and (iii) reporting of worst-group and macro-averaged metrics in the multi-class setting.

### v4

@@ 1 Introduction
Melanoma is the deadliest form of skin cancer, with an estimated 325,000 new cases (174,000 men and 151,000 women) and 57,000 deaths (32,000 men and 25,000 women) reported worldwide in 2020 [arnold2022global]. By 2040, incidence is projected to increase by nearly 50% to 510,000 new cases, and mortality by 68% to 96,000 deaths. Early detection, however, dramatically improves survival. For example, a population-wide screening project in Schleswig-Holstein, Germany, reported a reduction in melanoma mortality of 47% for men and 49% for women [katalinic2012does]. Similarly, screening efforts in Australia decreased melanoma deaths by 59% [watts2021association]. These findings underline a crucial insight: despite rising incidence, early and accessible screening can substantially reduce mortality.
Recent advances in Artificial Intelligence (AI), together with the widespread availability of high-resolution smartphone cameras, have created new opportunities for accessible melanoma pre-screening outside clinical environments. AI-assisted diagnostic systems can support clinicians during routine examinations and also provide human-understandable explanations for their decisions [lucieri2022exaid]. In principle, such systems could even enable remote pre-screening, allowing individuals to capture dermoscopic images themselves and receive preliminary risk assessments [chao2017smartphone, rat2018use]. These innovations have considerable societal potential, including earlier detection, reduced diagnostic delays, lower healthcare costs, and decreased clinician workload.
Figure 1: Real melanoma test data (left) such as MILK10k [MILK10k_2025] shows limited and imbalanced demographic coverage, while synthetic data (right) can provide complete combinations with balanced sampling, enabling more reliable fairness assessment.
At the same time, the deployment of AI in medicine requires a high level of trustworthiness, especially in light of regulatory developments such as the EU AI Act [european2019ethics]. Fairness and transparency in both datasets and trained models are essential conditions for reliable clinical use [bevan2022detecting, ansari2024algorithmic, aayushman2024fair, bie2024xcoop, pahde2023reveal, dembinsky2025unifying]. Fairness assessments depend on access to demographically well-characterized data. The International Skin Imaging Collaboration (ISIC) archive [gutman2016skin] provides one of the most comprehensive publicly available skin lesion datasets [gutman2016skin, codella2018skin, tschandl2018ham10000, codella2019skin, rotemberg2021patient, hernandez2024bcn20000], including metadata such as sex, age, and Fitzpatrick skin type [sachdeva2009fitzpatrick]. Despite its scale and quality, demographic imbalance remains a major limitation and complicates fairness evaluations.
@@
DeepGuide ([mallya2022deep]) DenseNet Derm7pt 0.788 (AUC)
MelaNet ([zunair2020melanoma]) VGG-GAP ISIC 2016 0.811 (AUROC)
Patch-Lesion ([gessert2020skin]) Dense121 HAM10000 0.855 (F1-Score)
SkinLesionDensenet ([panda2021comparative]) DenseNet201 ISIC 0.878 (F1-Score)
EFFNet ([ma2023effnet]) EfficientNetV2 HAM10000 0.932 (F1-Score)
sabir2024classification EfficientNet-B0 HAM10000 0.970 (F1-Score)
@@
Meanwhile, recent progress in Generative AI (GenAI), particularly diffusion models [ho2020denoising, rombach2022high], has enabled high-fidelity and attribute-controlled image synthesis. In this work, we explore whether these advances can support fairness evaluation in melanoma detection. Using the ISIC dataset and LightningDiT [yao2025reconstruction], a latent diffusion transformer [peebles2023scalable], we train a generative model capable of synthesizing realistic dermoscopic images conditioned on key patient attributes. We then examine whether such synthetic cohorts can serve as a reliable resource for assessing algorithmic fairness in publicly available melanoma classifiers. In summary, our work makes the following key contributions.
@@ 2.1 Generative Models for Skin Lesions
Synthetic data generation is widely used to address data scarcity, class imbalance, and privacy concerns in skin lesion datasets. Early approaches based on Generative Adversarial Networks (GANs) [goodfellow2020generative], such as MelanoGANs [baur2018melanogans], DCGAN [radford2015unsupervised], and LAPGAN [denton2015deep], demonstrated that realistic dermoscopic images can be synthesized even from limited data. DermGAN [ghorbani2019dermgan] and StyleGAN-based methods [bissoto2019skin, karras2019style] introduced finer control over lesion characteristics and improved downstream classifier performance. Later work, including StyleGAN2 [karras2020training] and federated setups [carrasco2022assessing], further improved stability, fidelity, and privacy. Diffusion models now dominate high-fidelity medical image synthesis. Stable Diffusion [rombach2022high] and DreamBooth [ruiz2023dreambooth], as applied in Cancer-Net SCa-Synth [tai2024cancer], support few-shot personalization and balanced dataset construction. Extensions of latent diffusion models have also improved control over skin type variation and helped address fairness gaps [wang2024majority]. Our work differs by focusing on fairness evaluation rather than augmentation. We train a state-of-the-art latent diffusion transformer based on LightningDiT [yao2025reconstruction] to generate attribute-controlled dermoscopic image cohorts. The model leverages all available ISIC metadata and allows systematic control over demographic and clinical attributes for constructing balanced intersectional test sets.
@@ 2.2 Melanoma Detection Models
Prior research on melanoma classification has employed a variety of datasets, including ISIC, Derm7pt [kawahara2018seven], and HAM [tschandl2018ham10000], the latter containing 10,015 dermoscopic images covering seven lesion types. These datasets differ in composition, imaging style, and demographic distribution, which complicates direct model comparison. A further challenge arises from the heterogeneous evaluation metrics reported in the literature. Some studies rely on the F1-score, whereas others report AUC or AUROC, making cross-model comparison nontrivial. Pretrained model availability also varies. Public weights exist for DeepGuide [mallya2022deep], MelaNet [zunair2020melanoma], and SkinLesionDensenet [panda2021comparative], but not for Patch-Lesion [gessert2020skin] or EFFNet [ma2023effnet]. In this work, we use publicly available pretrained models to ensure reproducibility and consistent evaluation.
@@ 2.3 Fairness in Skin Lesion Classification
Fairness evaluation has gained increasing attention due to persistent performance disparities across demographic subgroups. In [ansari2024algorithmic], the authors proposed an augmentation strategy that combines a source skin tone image with a diagnostic image to mitigate imbalance and skin tone bias. Other approaches incorporate fairness directly into model design. BiaslessNAS [sheng2024data] integrates fairness constraints throughout the neural architecture search process and reports improvements of 2.55% in accuracy and 65.50% in fairness over baseline models. Compared to previous augmentation-based fairness approaches, our method generates attribute-controlled synthetic images through text prompts. We evaluate fairness in existing pretrained melanoma classifiers [mallya2022deep, zunair2020melanoma, panda2021comparative] using balanced synthetic cohorts, allowing us to investigate the potential of prompt-driven generative models as a tool for systematic fairness assessment.
@@ 3.1 Synthetic Test Data Generation
Figure 2: Overall pipeline of our fairness testing workflow. We first train a generative model based on LightningDiT [yao2025reconstruction] using the ISIC dataset. Synthetic melanoma test images are then generated via systematic prompt templates. Finally, the synthetic images are used to assess fairness as measured via TPR difference in pre-trained melanoma detection models.
Data: Only publicly available data are used and no new patient information is collected. All dermoscopic images from the ISIC archive (2016–2024) are processed at a resolution of $256\times 256$ and encoded into the latent space using the VA-VAE, resulting in roughly 500,000 latent-image pairs. Each image is paired with a text description constructed from the available annotations, including sex, age group, lesion size, Fitzpatrick skin type, and diagnosis. Data with insufficient attributes are excluded. Text embeddings are produced using CLIP [radford2021learning], a contrastively trained vision-language model that embeds images and text into a shared latent space and provides strong semantic grounding for conditioning.
@@ 3.3 Fairness Assessment
Metric Definition: Since our synthetic test set contains only melanoma (i.e. positive) cases, we quantify subgroup disparities using TPR disparity [hardt2016equality], defined as the difference between the maximum and minimum true positive rates across PII groups. In a positive-only setting, the proportion of predicted positives in each group coincides with sensitivity. For a single decision threshold $\tau$ shared across all groups:
@@
We compute $\Delta_{\mathrm{TPR}}$ on synthetic cohorts stratified by sex, age group, and Fitzpatrick skin type. To generate a strictly positive-only test set, we use “melanoma” as the disease value in our systematic prompt configuration. The resulting measure, $\Delta_{\mathrm{TPR}}$, therefore provides a meaningful and statistically robust indicator of whether a model systematically favors or disadvantages particular demographic groups [stanley2022disproportionate, li2021estimating, kinyanjui2020fairness].
@@
Models: We evaluate three publicly available pretrained melanoma classifiers: DeepGuide [mallya2022deep], MelaNet [zunair2020melanoma], and SkinLesionDensenet [panda2021comparative]. These models were selected due to the availability of pretrained weights and their relevance in prior benchmark studies. MelaNet and SkinLesionDensenet are trained on ISIC, whereas DeepGuide is trained on the HAM10000 (HAM) dataset, providing complementary training distributions. Architecturally, DeepGuide and SkinLesionDensenet rely on DenseNet backbones, while MelaNet is based on a VGG-GAP design. This diversity in both training data and network architecture offers a representative foundation for examining fairness behavior across classifier families.
@@ 4.1 Generated Synthetic Skin Lesion Images
While the generated images are highly realistic, one limitation is immediately noticeable. The generated images do not became increasingly darker with higher Fitzpatrick skin type. For example, the generated images do not become progressively darker with higher Fitzpatrick skin types. Although this might be the intuitive expectation, prior work [tong2020investigating, kalb2023revisiting] shows that lighting, camera settings, and lesion visibility introduce substantial variability, making pixel intensity an unreliable indicator. Further details and visual comparisons with real data exhibiting the same issue are provided in the appendix.
@@ 4.2 Evaluation of Fairness Performance
Effect of Test Set Size: Figure 5 shows cumulative TPR as a function of the number of test images. The data count corresponds to the number of images used and illustrates how recognition accuracy changes as more test data are included. As a rule of thumb, as the number of test images increases, convergence to a stable accuracy is expected [rolnick2017deep]. Therefore, if accuracy trends for real and synthetic images converge, fairness simulations can be reliably performed using generated images. For sex, both male and female curves exhibit this convergence behavior, and synthetic testing closely tracks real-image performance as the available data volume increases. In contrast, the age-related curves for real and synthetic data do not fully converge, suggesting that the generator and real data distributions differ more strongly along the age dimension. Age is a factor with substantial intra-group variability (e.g., early versus late decades), which may contribute to these discrepancies and should be examined in future work.
@@ 5.1 Enhancing Dataset Use for Training
In this work, we generated synthetic image test data by training a custom LightningDiT [yao2025reconstruction]. According to Figure 3, the images show high realism, and Figure 5 confirms a similar trend in the cumulative TPR comparison between real and synthetic images. However, according to Figure 5, age TPR conversion differs between real and synthetic images. This may be due to the training procedure, which determines the precision of the label used for training. The next step to improve the synthetic image generation model is to include a detail attribute in the training procedure of our generative model. Another direction is to add multiple datasets and use different image capturing devices to make the dataset and thus generation process more robust.
@@ 5.2 Quality Control of the Synthetic Images
In our work, we evaluate fairness performance through 81,600 synthetic images generated in a pattern presented in Section 3.2. To ensure the quality of the generated images, further quality control models could be used that evaluate the reliability of the images [pmlr-v136-kwon20a]. The approach may, for example, be to implement quality-judging models for each age, sex, and skin color. Also, due to the significant recent growth of visual language models (VLMs) [zhang2024vision], VLM-as-a-judge may be an approach to perform quality evaluation before model prediction. Hence, the next step in the current architecture is also to consider new models for quality assurance of the test images.
@@ 5.3 Understanding Causes of Unfairness
Our results verified that synthetic image generation can serve as test data for evaluating the fairness of pre-trained models. For instance, when examining the performance of pre-trained models, we confirm that for sex and skin type, MelaNet shows the most fair behavior, with $\Delta_{\mathrm{TPR}}$ of 0.0115 and 0.1322. For age SkinLesionDensnet performs the best with 0.0682. Although we have managed to measure $\Delta_{\mathrm{TPR}}$ for model fairness comparison, we still lack a method for discovering the causes or factors of unfairness. Hence, the next step is to apply concept-based fairness [nejadgholi2022towards, pendyala2022concept, tong2020investigating] evaluation to understand the actual cause of the fairness issue.
@@ 5.4 Various Diagnoses for Testing
We focused on TPR assessed through the disparity in true positive rates ($\Delta_{\mathrm{TPR}}$) under a melanoma-only (positive) test design. To broaden the scope of fairness evaluation, we aim to extend the generator and evaluation protocol to multiple diagnostic categories (e.g., benign nevi and other lesion types) and to multi-class classifiers in the future. This will enable (i) analysis of misclassification patterns (false positives and false negatives) by PII group, (ii) computation of additional group fairness criteria beyond $\Delta_{\mathrm{TPR}}$, such as equalized odds [hardt2016equality], and (iii) reporting of worst-group and macro-averaged metrics in the multi-class setting.